OpenStudy (anonymous):

A circular table has the DIAMETER OF 48". The table can be enlarged by insterting two rectangular leaves each 18"wide and 48" long. What is the AREA of the expanded table? Round to nearest tenth....HELP!!! :)

OpenStudy (campbell_st):

won't the new area be Area of the circle + area of the rectangles that are added for the circle you need to radius. The question only gives a diameter so the total area is \[Area = \pi r^2 + 2 \times length \times width\] hope this helps.

OpenStudy (anonymous):

im getting to big of a number, im getting 3536.64

OpenStudy (anonymous):



OpenStudy (campbell_st):

very close to my answer... did you use 3.14 for pi...?

OpenStudy (anonymous):

yes I did 24x24x3.14= 1808.644 then I did 18x48=864

OpenStudy (anonymous):

is the answer 3536.64"

OpenStudy (campbell_st):

ok.. thats fine I used pi on my calculator... and there is a difference of 0.9. So well done